Calculate the final pressure, in atmospheres, for each of the following, if volume and amount of gas do not change.

You may want to reference (Pages 286 - 288) Section 8.4 while completing this problem.

Part A

Part complete

A gas with an initial pressure of 1.26 atm at 65 ∘C is cooled to − 25 ∘C .

Express the pressure to three significant figures and include the appropriate units.

--

SubmitPrevious AnswersRequest Answer

Part B

Part complete

A sample of N2 with an initial pressure of 780. mmHg at − 81 ∘C is heated to 26 ∘C .

Express the pressure to three significant figures and include the appropriate units.

--

P2

P 2

=
P2

P 2
Answer #1

A)

Given:

Pi = 1.26 atm

Ti = 65.0 oC

= (65.0+273) K

= 338 K

Tf = -25.0 oC

= (-25.0+273) K

= 248 K

use:

Pi/Ti = Pf/Tf

1.26 atm / 338 K = Pf / 248 K

Pf = 0.924 atm

Answer: 0.924 atm

B)

Given:

Pi = 780 mmHg

Ti = -81.0 oC

= (-81.0+273) K

= 192 K

Tf = 26.0 oC

= (26.0+273) K

= 299 K

use:

Pi/Ti = Pf/Tf

780 mmHg / 192 K = Pf / 299 K

Pf = 1215 mmHg

= 1215 / 760 atm

= 1.60 atm

Answer: 1.60 atm